Nihal is borrowing Rs. 50,000 to buy ABC assets. If he pays equal installments for 25


years and 4% interest on the outstanding balance, what is the amount of installment?


What shall be the amount of installment if quarterly payments are required to be made?

1
Expert's answer
2021-10-31T18:28:20-0400

Let's apply the formula for the annuity payment on the loan:

n=25

i=0.04

"x=S(i+\frac{i}{(1+i)^n-1}=50000(0.04+\frac{0.04}{(1+0.04)^{25}-1}=3200.60"

if quarterly payments: 

"n=4\times25=100"

"i=\frac{0.04}{4}=0.01"

"x=S(i+\frac{i}{(1+i)^n-1}=50000(0.01+\frac{0.01}{(1+0.01)^{100}-1}=793.29"